OpenStudy (anonymous):

\[\sqrt[3]{28x}=-4\]

OpenStudy (anonymous):

\[\sqrt[3]{28x}=-4\]

OpenStudy (ybarrap):

$$ \large{ \sqrt[3]{28x}=-4\\ 28x=(-4)^{1/3}\\ x=\cfrac{(-4)^{1/3}}{28}\\ =2^{2/3}\times\sqrt[3]{-1}~ } $$ Does that help?

OpenStudy (anonymous):

Not really @ybarrap